Will you please help me with this problem?
Write a system of two equations in two unknowns. Solve the system by using the addition method. « Mike paid $2.98 for three doughnuts and two coffees. The next day, he paid $1.64 for two doughnuts and one coffee. How much does it cost for one doughnut and one coffee?
Thank you!

Answer by ptaylor(1894)


Let x=cost of 1 doughnut
And y=cost of 1 coffee
Now we are told that mike paid $2.98 for 3 doughnuts and 2 coffees
so 3x+2y=$2.98
Next day we are told that Mike paid $1.64 for 2 doughnuts and 1 coffee
That would be 2x+y=$1.64
Now we have our two equations:
(1) 3x+2y=$2.98
(2) 2x+y=$1.64
Multiply Equation #(2) by (-2) and we get:
(2a) -4x-2y=-$3.28
(1) 3x+2y=$2.98
Add (1) and (2a) and we get:
-x=-$.30
x=30 cents cost of 1 doughnut
substitute in (1) and we have:
3($.30)+2y=$2.98
$.90+2y=$2.98 subtract $.90 from both sides:
2y=$2.98-$.90
2y=$2.08
y=$1.04 cost of 1 coffee
ck substitute in (1)
3($.30)+2($1.04)=$2.98
$.90+$2.08=$2.98
$2.98=$2.98
Substitute in (2)
2($.30)+$1.04=$1.64
$.60+$1.04=$1.64
$1.64=$1.64

Hope this helps. Happy holidays.----ptaylor